# FAIR-TAT: Improving Model Fairness Using Targeted Adversarial Training

### v1

@@ 1 Introduction
Deep neural networks have led to tremendous success in a wide range of computer vision applications [13, 9, 16]. One of the most relevant remaining challenges is their lack of robustness towards small domain shifts [24], common corruptions [14, 15] and adversarially perturbed examples [10, 18, 8]. This limits their usage in safety-critical applications [21]. Adversarial examples are artificially crafted samples, which do not alter the semantic meaning of natural samples. However, they add a small amount of noise that alters the model’s prediction. As a prominent defense mechanism to such adversaries, models are trained using adversarial training (AT) algorithms [44, 23, 34, 41], where adversarial samples along with clean samples are utilized during model training to build a more robust classifier. The most commonly used adversarial examples during AT are generated through Fast Gradient Sign Method (FGSM) [10] and Projected Gradient Descent (PGD) [23]. These methods are overall efficient to compute and have shown to improve the overall model robustness [34, 23, 41].
@@ 2 Related Work
Adversarial Fairness Problem.
Adversarial Fairness in machine learning is a growing concern, as models often favor specific groups of classes while underperforming on others after being trained adversarially. This discrimination undermines the reliability of machine learning models. Recent studies have investigated class biases in terms of model accuracy and robustness, advocating for fair training processes that classify all classes equally well [36, 31, 37]. These works also show that while AT can improve overall robustness, it often ignores discrepancies in the robustness of individual classes, amplifying unbalance in class-wise robustness. Data distribution and adversarial learning algorithms can contribute to these class-wise discrimination problems, making hard classes more vulnerable to attacks.
However, there exists classes which exhibit higher robustness performance but also frequently tend to be confused with other classes. Therefore, in this work, we focus on customizing the training configuration of Adversarial Training (AT) based on class-wise confusion rates, which are evaluated using class-wise false positive scores.
@@ 4 Empirical Analysis on $C_{FPS}$
We conduct an experiment by evaluating $C_{FPS}$ on the CIFAR-10 dataset, employing robust models trained adversarially that are sourced from a standardized adversarial robustness benchmark [5, 12]. Figure 1 illustrates the $C_{FPS}$ scores for CIFAR-10 classes. A clear trend emerges in the $C_{FPS}$ scores across different adversarially trained architectures. The horizontal axis in Figure 1 represent the 10 classes of CIFAR-10 in order. Notably, class "cat", is frequently wrongly assigned to samples from other classes. Similarly, class "airplane", although considered one of the easier classes due to its relatively high robust accuracy, also shows significant misclassification rates. This indicates a need for a deeper understanding of the decision boundaries of these classes when perturbed.
To address the high confusion rates of specific classes, our work uses $C_{FPS}$ to model a prior multinomial distribution of classes. This allows us to sample highly confused classes more frequently during targeted adversarial training, where we employ targeted adversaries towards these classes to harden the model against class specific perturbations.
@@ 6 Experiments
Table 1: Comparison of PGD evaluations of FAIR-TAT framework with other methods focused on model fairness using PRN-18 on CIFAR-10 and CIFAR-100. Our method is marked with ◆.
@@
Table 2: Comparison of AutoAttack and Squares evaluations of FAIR-TAT framework with other methods focused on model fairness using PRN-18 on CIFAR-10 and CIFAR-100. Our method is marked with ◆.
@@
Corruption Type FAIR_TAT FAIR_TAT (EMA) FAIR_TAT (CFA) FAIR_TAT (T) FAIR_TAT (T+EMA) FAIR_TAT (T+CFA) FAIR_TAT (F) FAIR_TAT (F+EMA) FAIR_TAT (F+CFA)

### v2

@@ 1 Introduction
Deep neural networks have led to tremendous success in a wide range of computer vision applications [19, 13, 23]. One of the most relevant remaining challenges is their lack of robustness towards small domain shifts [33], common corruptions [20, 22] and adversarially perturbed examples [14, 25, 11, 3]. This limits their usage in safety-critical applications [29]. Adversarial examples are artificially crafted samples, which do not alter the semantic meaning of natural samples. However, they add a small amount of noise that alters the model’s prediction. As a prominent defense mechanism to such adversaries, models are trained using adversarial training (AT) algorithms [54, 31, 44, 51], where adversarial samples along with clean samples are utilized during model training to build a more robust classifier. The most commonly used adversarial examples during AT are generated through Fast Gradient Sign Method (FGSM) [14] and Projected Gradient Descent (PGD) [31]. These methods are overall efficient to compute and have shown to improve the overall model robustness [44, 31, 51].
@@ 2 Related Work
Adversarial Robustness.
Adversarial robustness is a crucial concept for assessing the resilience of deep neural networks against adversarial attacks [14, 25, 11, 3]. Significant progress has been made in evaluating robustness across diverse adversarially trained model architectures thoroughly [16, 8], while some studies have aimed to unify research efforts on architectural design choices and their influence on robustness [21]. Furthermore, advancements in adversarial robustness have been achieved across tasks such as image classification, restoration, and segmentation by enhancing or redesigning robust architectures [18, 2, 1, 17, 28, 21]. Despite these efforts, which focus on improving overall robust accuracy, consistent variations in robustness across specific classes have been observed. These findings highlight the nuanced challenges in achieving uniformly robust performance across all categories.
@@
Adversarial Fairness Problem.
Adversarial Fairness in machine learning is a growing concern, as models often favor specific groups of classes while underperforming on others after being trained adversarially. This discrimination undermines the reliability of machine learning models. Recent studies have investigated class biases in terms of model accuracy and robustness, advocating for fair training processes that classify all classes equally well [46, 41, 47, 32]. These works also show that while AT can improve overall robustness, it often ignores discrepancies in the robustness of individual classes, amplifying unbalance in class-wise robustness. Data distribution and adversarial learning algorithms can contribute to these class-wise discrimination problems, making hard classes more vulnerable to attacks.
However, there exist classes which exhibit higher robustness performance but also frequently tend to be confused with other classes. Therefore, in this work, we focus on customizing the training configuration of Adversarial Training (AT) based on class-wise confusion rates, which are evaluated using class-wise false positive scores.
@@ 4 Empirical Analysis on $C_{FPS}$
We conduct an experiment by evaluating $C_{FPS}$ on the CIFAR-10 dataset, employing robust models trained adversarially that are sourced from a standardized adversarial robustness benchmark [8, 16] and a foundation model, DINOv2[35]. Figure 1 illustrates the $C_{FPS}$ scores for CIFAR-10 classes. A clear trend emerges in the $C_{FPS}$ scores across different adversarially trained architectures. The horizontal axis in Figure 1 represent the 10 classes of CIFAR-10 in order. Notably, class "C4: cat", is frequently wrongly assigned to samples from other classes. Similarly, class "C1: airplane", although considered one of the easier classes due to its relatively high robust accuracy, also shows significant misclassification rates. This indicates a need for a deeper understanding of the decision boundaries of these classes when perturbed. To address the high confusion rates of specific classes, our work uses $C_{FPS}$ to model a prior multinomial distribution of classes. This allows us to sample highly confused classes more frequently during targeted adversarial training, where we employ targeted adversaries towards these classes to harden the model against class specific perturbations.
@@ 6 Experiments
Table 1: Comparison of PGD evaluations of FAIR-TAT framework with other methods focused on model fairness using PRN-18 on CIFAR-10 and CIFAR-100 with $\epsilon=8/255$. Our method is marked with ◆.
@@
Table 2: Comparison of AutoAttack and Squares evaluations of FAIR-TAT framework with other methods focused on model fairness using PRN-18 on CIFAR-10 and CIFAR-100 with $\epsilon=8/255$. Our method is marked with ◆.
@@
Corruption Type (C) FAIRTAT FAIRTAT (EMA) FAIRTAT (CFA) FAIRTAT (T) FAIRTAT (T+EMA) FAIRTAT (T+CFA) FAIRTAT (F) FAIRTAT (F+EMA) FAIRTAT (F+CFA)
@@ 7 Conclusion
Acknowledgements
Steffen Jung and Margret Keuper acknowledge funding by the DFG Research Unit 5336 - Learning to Sense.
Supplementary Material
In this document, we provide the additional details and experimental results of our approach. The supplementary material is structured as follows:
(A) Ablation Study
  • Fairness in adversarially trained models.
  • Ablations on uniform sampling of targets.
  • Ablations on constant perturbation margin during AT.
  • Ablations on epsilon scaling.
(B) Additional Experimental Results
  • Experimental results with PRN-18 on CIFAR-10 over mutliple seeds.
  • Experimental results with XCiT-S12 on CIFAR-10.
  • Overall accuracy results on common corruptions.
Appendix A Ablation Study
A.1 Fairness in Adversarially Trained Models.
Figure 3: Class-wise accuracy $C_{acc}$ on the CIFAR-10 dataset using adversarially trained models.
Table 5: Comparison of PGD evaluations of FAIR-TAT framework with its variants. ($\epsilon=8/255$)
Method PRN-18 (Clean Accuracy) PRN-18 (Robust Accuracy)
Overall Worst Class Overall Worst Class
FAIR-TAT (UTS) $85.0\pm 0.2$ $70.9\pm 2.3$ $46.2\pm 0.5$ $18.4\pm 0.9$
FAIR-TAT (CTS) ◆ $\mathbf{85.5\pm 0.3}$ $\mathbf{71.3\pm 1.8}$ $\mathbf{46.6\pm 0.6}$ $\mathbf{19.7\pm 1.1}$
FAIR-TAT (UTS) + EMA $85.1\pm 0.6$ $72.2\pm 0.8$ $47.4\pm 0.3$ $21.2\pm 0.5$
FAIR-TAT (CTS) + EMA ◆ $\uline{\mathbf{86.0\pm 0.4}}$ $\uline{\mathbf{73.0\pm 0.7}}$ $\mathbf{47.7\pm 0.3}$ $\mathbf{22.1\pm 0.9}$
FAIR-TAT (UTS) + CFA $83.0\pm 0.4$ $70.2\pm 1.3$ $47.8\pm 0.2$ $23.3\pm 0.8$
FAIR-TAT (CTS) + CFA ◆ $\mathbf{84.8\pm 0.3}$ $\mathbf{72.0\pm 0.9}$ $\uline{\mathbf{48.3\pm 0.5}}$ $\uline{\mathbf{24.6\pm 1.0}}$
Table 6: Comparison of PGD evaluations of FAIR-TAT framework with constant perturbation margin during training. ($\epsilon=8/255$)
Method PRN-18 (Clean Accuracy) PRN-18 (Robust Accuracy)
Overall Worst Class Overall Worst Class
AT $84.0\pm 0.2$ $66.4\pm 1.3$ $\mathbf{45.6\pm 0.2}$ $16.9\pm 1.1$
FAIR-TAT (UTS) $87.3\pm 0.1$ $\mathbf{73.8\pm 2.9}$ $43.6\pm 0.7$ $\mathbf{19.0\pm 1.9}$
FAIR-TAT (CTS) $\mathbf{87.3\pm 0.3}$ $73.2\pm 1.7$ $43.5\pm 0.6$ $17.7\pm 2.7$
AT + EMA $84.7\pm 0.3$ $68.5\pm 0.9$ $\mathbf{47.6\pm 0.1}$ $18.6\pm 0.6$
FAIR-TAT (UTS) + EMA $\uline{\mathbf{88.0\pm 1.5}}$ $\uline{\mathbf{75.9\pm 1.0}}$ $44.5\pm 0.5$ $\mathbf{20.1\pm 0.4}$
FAIR-TAT (CTS) + EMA $87.8\pm 0.2$ $74.9\pm 0.9$ $44.6\pm 0.3$ $19.0\pm 1.2$
AT + CFA $84.6\pm 0.3$ $69.5\pm 1.5$ $\uline{\mathbf{48.3\pm 0.2}}$ $21.8\pm 0.6$
FAIR-TAT (UTS) + CFA $\mathbf{87.3\pm 0.4}$ $\mathbf{74.8\pm 1.1}$ $45.9\pm 0.5$ $\uline{\mathbf{23.0\pm 2.4}}$
FAIR-TAT (CTS) + CFA $86.0\pm 3.2$ $73.3\pm 7.4$ $44.8\pm 0.3$ $22.3\pm 1.4$
We conducted an experiment on the CIFAR-10 dataset using robust models trained adversarially, sourced from a standard adversarial benchmark [8] and additional foundation model, DINOv2[35] trained adversarially. Figure 3 shows the class-wise robustness accuracies ($C_{acc}$) on the CIFAR-10 dataset. The ten classes of CIFAR-10, namely ’airplane’, ’automobile’, ’bird’, ’cat’, ’deer’, ’dog’, ’frog’, ’horse’, ’ship’, and ’truck’, are denoted as C1 to C10 respectively. The Figure 3 illustrates the class-wise accuracy of different adversarially trained models on clean validation samples from the CIFAR-10 dataset. The disparity in class-wise accuracies are clearly visible, which means the adversarially trained models are not fair. The dotted lines in the Figure 3 represent the overall accuracies of the respective models. The overall accuracy of each model allows us to categorize classes into two groups: Easy classes and Hard classes. This categorization is based on whether a class has a class-wise accuracy above or below the model’s average overall accuracy. Easy classes are easy to predict and hard classes are often hard to distinguish. This pattern remains consistent across different model architectures. Therefore training a fair robust model is very important. Our approach aims to work on this robust fairness issue.
A.2 Ablation on Uniform Sampling of Targets
In this section, we compare our approach, FAIR-TAT, which samples targets based on the class-wise false positive prior distribution, to a uniform target sampling method during targeted adversarial training. This comparison demonstrates the effectiveness of using class-wise false positive information in target sampling, highlighting its benefits in preserving overall accuracy and improving the worst-class accuracy.
We refer to our approach, which employs class-wise false positive target selection, as FAIR-TAT(CTP). Conversely, we denote the approach that uses uniform target sampling during training as FAIR-TAT(UTP) in Table 5. The results in the table present PGD evaluations using PRN-18 on the CIFAR-10 dataset. It should also be noted that the perturbation margins for the classes are customized as described in [46] during the training of these approaches. From Table 5, it is clear that sampling targets based on class-wise false positive scores improves model robustness and fairness compared to uniform sampling of targets. This improvement arises from assigning more weight to classes that are more prone to misclassification during training. Thus, our results demonstrate the effectiveness of incorporating class-wise false positive scores when sampling targets for targeted adversarial training scenario of our approach. We observe the same trend with different weight averaging techniques EMA and CFA as well.
A.3 Ablations on Constant Perturbation Margin during AT.
In this section, we assess the efficacy of our approach when the perturbation margin $\epsilon$ is kept constant during the training process. For a fair comparison, we use vanilla adversarial training (AT) as a baseline, which also employs a constant perturbation margin on the CIFAR-10 dataset using the PRN-18 architecture. Additionally, we evaluate our method under different target sampling schemes, as well as conventional AT, using various weight averaging schemes.
Table 6 presents the results of our approach under two conditions: Uniform Target Sampling (UTS) and Class-wise False Positive Target Sampling (CTS), compared to the baseline AT. The results indicate that FAIR-TAT (UTS) with a constant perturbation margin $\epsilon$ performs well in terms of fairness, showing improved worst-class accuracies compared to the baseline. Furthermore, FAIR-TAT (UTS) and FAIR-TAT (CTS) increases the overall clean accuracy, although the overall robustness decreases significantly when compared to the baseline. This observation holds across different weight averaging schemes.
For results using a customized perturbation margin, refer to Table 5. From Table 5, it is evident that adjusting the perturbation margin during targeted adversarial training allows for adjusting attack strengths across different classes. Harder classes are assigned a smaller margin, while easier classes receive a larger margin, as described in [46]. Interestingly, FAIR-TAT with uniform target sampling performs slightly better with a constant $\epsilon$ than class-wise false positive target sampling with a constant $\epsilon$. FAIR-TAT (UTS) also demonstrates superior fairness in terms of clean sample accuracy when considering all versions of our method.
However, the combination of customized margins and class-wise false positive target sampling (CTP) in FAIR-TAT offers better trade-offs between robustness and fairness, considering both robust and clean samples. Thus, comparing Table 5 and Table 6, we conclude that the combination of FAIR-TAT (CTP), customized perturbation margins, and effective weight averaging leads to a robust and fair classifier.
Figure 4: Epsilon scaling during AT for FAIR-TAT and CFA method respectively on 10 classes of CIFAR10 in order.
Table 7: Comparison of PGD evaluations of FAIR-TAT framework with other methods focused on model fairness using PRN-18 on CIFAR-10 over 5 seeds with $\epsilon=8/255$. Our method is marked with ◆.
Method PRN-18 (Clean Accuracy) PRN-18 (Robust Accuracy)
Overall Worst Class Overall Worst Class
AT $83.5\pm 0.3$ $66.8\pm 2.7$ $\mathbf{48.2\pm 0.3}$ $19.6\pm 2.0$
FAIR-TAT ◆ $\mathbf{85.5\pm 0.2}$ $\mathbf{71.3\pm 2.1}$ $46.6\pm 0.2$ $\mathbf{19.7\pm 2.0}$
AT + EMA $83.8\pm 0.2$ $66.6\pm 1.3$ $\mathbf{49.6\pm 0.2}$ $21.0\pm 0.5$
FAIR-TAT + EMA ◆ $\mathbf{86.0\pm 0.2}$ $\mathbf{73.0\pm 0.5}$ $47.7\pm 0.2$ $\mathbf{21.7\pm 0.5}$
AT + CFA $83.8\pm 0.2$ $68.1\pm 1.1$ $\mathbf{50.1\pm 0.2}$ $22.8\pm 1.4$
FAIR-TAT + CFA◆ $\mathbf{84.8\pm 1.1}$ $\mathbf{72.0\pm 3.4}$ $48.3\pm 1.1$ $\mathbf{24.6\pm 2.4}$
TRADES $\mathbf{82.0\pm 0.4}$ $64.7\pm 1.4$ $52.9\pm 0.4$ $\mathbf{25.9\pm 1.7}$
FAIR-TAT + TRADES ◆ $81.7\pm 0.1$ $\mathbf{67.7\pm 1.1}$ $\mathbf{52.3\pm 1.1}$ $25.5\pm 0.7$
TRADES + EMA $\mathbf{82.4\pm 0.1}$ $65.3\pm 0.7$ $\uline{\mathbf{53.8\pm 0.1}}$ $25.3\pm 0.6$
FAIR-TAT + TRADES + EMA ◆ $82.1\pm 0.2$ $\mathbf{68.4\pm 0.8}$ $53.1\pm 0.2$ $\mathbf{26.8\pm 0.8}$
TRADES + CFA $\mathbf{82.3\pm 0.2}$ $65.4\pm 0.7$ $\uline{\mathbf{53.7\pm 0.2}}$ $25.2\pm 0.2$
FAIR-TAT + TRADES + CFA ◆ $82.1\pm 0.2$ $\mathbf{68.6\pm 1.0}$ $53.1\pm 0.2$ $\uline{\mathbf{26.9\pm 0.8}}$
FAT $84.8\pm 0.3$ $69.3\pm 1.1$ $\mathbf{48.0\pm 0.3}$ $18.7\pm 0.8$
FAIR-TAT + FAT ◆ $\mathbf{86.6\pm 0.3}$ $\uline{\mathbf{77.2\pm 0.9}}$ $46.1\pm 0.3$ $\mathbf{22.7\pm 0.8}$
FAT + EMA $85.1\pm 0.2$ $69.5\pm 1.4$ $\mathbf{49.0\pm 0.2}$ $19.5\pm 0.9$
FAIR-TAT + FAT+ EMA ◆ $\uline{\mathbf{86.8\pm 0.2}}$ $\mathbf{75.4\pm 1.4}$ $47.7\pm 0.2$ $\mathbf{22.3\pm 0.9}$
FAT + CFA $85.0\pm 0.6$ $71.0\pm 1.2$ $\mathbf{51.0\pm 0.6}$ $23.4\pm 0.3$
FAIR-TAT + FAT + CFA ◆ $\mathbf{86.1\pm 0.2}$ $\mathbf{74.9\pm 1.1}$ $48.2\pm 0.2$ $\mathbf{24.4\pm 5.2}$
FRL $82.8\pm 0.1$ $71.4\pm 2.4$ $45.7\pm 0.3$ $24.4\pm 1.0$
FRL + EMA $83.6\pm 0.3$ $69.5\pm 0.7$ $46.3\pm 0.2$ $24.8\pm 0.4$
BAT + TRADES $86.5\pm 0.1$ $73.4\pm 1.4$ $49.8\pm 0.2$ $22.1\pm 0.9$
WAT + TRADES $81.2\pm 0.3$ $65.9\pm 2.0$ $47.1\pm 0.3$ $26.3\pm 0.9$
Clean Training $94.0\pm 0.3$ $79.4\pm 2.4$ $2.7\pm 0.7$ 00
Table 8: Comparison of AutoAttack and Squares evaluations of FAIR-TAT framework with other methods focused on model fairness using PRN-18 on CIFAR-10 over 5 seeds with $\epsilon=8/255$. Our method is marked with ◆.
Method PRN-18 (AutoAttack) PRN-18 (Squares)
Overall Worst Class Overall Worst Class
AT $\mathbf{45.7\pm 0.3}$ $15.4\pm 1.6$ $51.3\pm 0.2$ $19.4\pm 1.3$
FAIR-TAT ◆ $45.0\pm 0.2$ $\mathbf{18.7\pm 1.7}$ $\mathbf{51.8\pm 0.1}$ $\mathbf{26.7\pm 0.7}$
AT + EMA $\mathbf{45.6\pm 0.2}$ $15.4\pm 1.8$ $51.7\pm 0.2$ $19.7\pm 0.4$
FAIR-TAT + EMA ◆ $45.0\pm 0.3$ $\mathbf{18.8\pm 0.4}$ $\mathbf{51.8\pm 0.3}$ $\mathbf{26.7\pm 0.7}$
AT + CFA $47.4\pm 0.8$ $19.3\pm 0.3$ $51.7\pm 0.2$ $19.9\pm 0.4$
FAIR-TAT+CFA ◆ $\mathbf{47.0\pm 1.3}$ $\uline{\mathbf{24.8\pm 2.1}}$ $\mathbf{52.6\pm 0.8}$ $\uline{\mathbf{31.3\pm 1.2}}$
TRADES $\mathbf{49.8\pm 0.2}$ $18.7\pm 1.1$ $\mathbf{54.0\pm 0.3}$ $22.4\pm 0.2$
FAIR-TAT + TRADES ◆ $48.6\pm 0.1$ $\mathbf{20.8\pm 1.3}$ $52.6\pm 0.2$ $\mathbf{26.1\pm 0.4}$
TRADES + EMA $\mathbf{49.8\pm 0.2}$ $18.6\pm 0.4$ $\mathbf{54.0\pm 0.3}$ $22.4\pm 0.2$
FAIR-TAT + TRADES + EMA ◆ $48.6\pm 0.3$ $\mathbf{21.1\pm 0.6}$ $52.6\pm 0.2$ $\mathbf{26.1\pm 0.4}$
TRADES + CFA $\uline{\mathbf{50.3\pm 0.1}}$ $21.3\pm 0.3$ $\uline{\mathbf{54.2\pm 0.3}}$ $22.6\pm 0.4$
FAIR-TAT + TRADES + CFA ◆ $49.8\pm 0.3$ $\mathbf{24.0\pm 0.8}$ $53.0\pm 0.5$ $\mathbf{25.8\pm 0.9}$
FAT $44.2\pm 0.3$ $16.0\pm 1.1$ $50.8\pm 0.2$ $20.8\pm 0.1$
FAIR-TAT + FAT ◆ $\mathbf{44.6\pm 0.2}$ $\mathbf{17.9\pm 0.8}$ $\mathbf{51.5\pm 0.4}$ $\mathbf{26.3\pm 0.3}$
FAT + EMA $44.2\pm 0.1$ $15.9\pm 0.9$ $50.8\pm 0.2$ $20.8\pm 0.1$
FAIR-TAT + FAT+ EMA ◆ $\mathbf{44.6\pm 0.2}$ $\mathbf{18.1\pm 0.7}$ $\mathbf{51.5\pm 0.4}$ $\mathbf{26.3\pm 0.3}$
FAT + CFA $\mathbf{49.4\pm 0.1}$ $22.6\pm 0.8$ $\mathbf{53.4\pm 0.1}$ $24.1\pm 0.5$
FAIR-TAT + FAT + CFA ◆ $44.3\pm 0.3$ $\mathbf{22.1\pm 2.1}$ $49.0\pm 0.6$ $\mathbf{24.4\pm 0.3}$
FRL $44.0\pm 0.2$ $23.2\pm 1.2$ $49.7\pm 0.3$ $24.6\pm 0.9$
FRL + EMA $44.2\pm 0.3$ $23.9\pm 0.4$ $50.8\pm 0.6$ $24.9\pm 0.3$
BAT + TRADES $45.9\pm 0.4$ $18.7\pm 1.2$ $52.3\pm 0.4$ $23.8\pm 0.7$
WAT + TRADES $47.1\pm 0.3$ $24.5\pm 1.1$ $51.7\pm 0.3$ $25.6\pm 0.9$
A.4 Ablations on Epsilon Scaling.
It is well known that targeted adversaries are weaker than untargeted adversaries. Thus, we adjust the perturbation margin for each class during adversarial training (AT) using the update equation:
$\epsilon_{k}\leftarrow(\lambda_{1}+r_{k})\epsilon.$
This perturbation margin update for each class, $\epsilon_{k}$, depends on the robust class-wise accuracy performance, $r_{k}$, so that the perturbation margin for the individual class adjusts according to the model’s performance on this class. This ensures the strength of the adversaries for each class during AT. The Figure 4 showcases the class-wise perturbation margin for both our approach FAIR-TAT and CFA (which uses an adaptable class-wise perturbation margin during untargeted adversarial training) on the CIFAR10 dataset using the PRN-18 architecture. This provides intuition regarding the strength of adversaries for both targeted and untargeted setups during the dynamic training scheme. The perturbation margin, $\epsilon$, for hard classes is lower than for easier classes, and the $\epsilon$ values for the targeted scenario are greater than those for the untargeted scenario during the training. This balance ensures that the attack strength is balanced considering the targeted and untargeted adversaries.
Furthermore, the variance introduced by including targeted adversaries in the adversarial training of our approach is enhanced when we sample the targets from the prior distribution of class-wise false positive scores. This additional degree of freedom in the training process is intentional and helps to guide the sampling towards vulnerable classes during adversarial training, ultimately improving model fairness.
Appendix B Additional Experimental Results
B.1 Experimental Results with PRN-18 on CIFAR-10 over Mutliple Seeds.
In this section, we present the evaluations of our approach using PRN-18 on CIFAR-10 over 5 random seeds. Table 7 highlights the PGD evaluations, comparing our method, adapted to other approaches, to baseline methods on CIFAR-10. From our observations, it is clear that our approach, utilizing the weight averaging technique from CFA [46], provides enhanced model fairness, as it incorporates a fairness-aware weight averaging scheme.
Similarly, Table 8 presents the evaluations using standard AutoAttack and a black box attack named Square Attack, comparing our method against baseline approaches on CIFAR-10 over 5 seeds. The results follow the same trend, with our approach demonstrating improved fairness through the appropriate weight averaging scheme, while maintaining overall robust accuracy. Also, our approach provides better performance considering robustness and fairness on clean samples.
Table 9: Overall accuracies of FAIR-TAT method on common corruptions along with combination of TRADES (T) and FAT(F) approaches using PreActResNet-18 on CIFAR-10C dataset.
Corruption Type (C) FAIRTAT FAIRTAT (EMA) FAIRTAT (CFA) FAIRTAT (T) FAIRTAT (T+EMA) FAIRTAT (T+CFA) FAIRTAT (F) FAIRTAT (F+EMA) FAIRTAT (F+CFA)
gaussian_noise 80.6 81.6 79.0 78.5 76.3 76.3 82.2 82.5 73.3
shot_noise 81.7 82.6 80.1 79.3 77.5 77.3 83.1 83.6 74.5
speckle_noise 81.5 82.6 80.2 79.3 77.3 77.3 82.9 83.6 74.1
impulse_noise 71.5 72.7 72.3 74.9 72.2 72.2 73.0 74.1 66.5
defocus_blur 80.3 80.6 76.5 76.8 76.6 76.6 81.0 81.2 68.1
gaussian_blur 78.0 78.3 73.7 74.6 74.5 74.4 78.5 78.7 64.4
glass_blur 76.3 76.4 71.9 73.2 73.1 73.0 76.9 77.2 62.4
motion_blur 79.1 79.5 75.2 75.6 75.7 75.7 80.2 80.5 65.5
zoom_blur 78.8 79.5 77.2 75.2 76.3 76.2 80.1 80.3 72.0
snow 56.1 56.4 50.6 53.4 53.0 53.0 58.3 57.7 39.6
frost 82.6 82.7 79.9 77.0 78.2 78.2 83.1 83.0 74.3
fog 41.0 40.5 35.6 38.4 38.1 38.0 42.5 41.6 28.7
brightness 79.2 79.8 75.9 76.0 75.6 75.5 80.4 80.6 67.1
contrast 83.4 83.9 80.8 80.0 79.6 79.5 84.3 84.5 74.1
elastic_transform 82.9 83.6 80.6 80.2 79.6 79.5 84.1 84.4 74.7
pixelate 80.0 81.0 79.0 78.3 76.8 76.8 81.7 81.9 73.6
jpeg_compression 81.6 82.3 80.6 78.6 77.6 77.6 82.6 83.0 76.0
defocus_blur 75.8 75.8 71.3 68.1 70.8 70.8 76.2 75.8 63.7
Table 10: Overall accuracies of baselines on common corruptions along with combination of TRADES (T) and FAT(F) approaches using PreActResNet-18 on CIFAR-10C dataset.
Corruption Type (C) AT AT (EMA) AT (CFA) AT (T) AT (T+EMA) AT (T+CFA) AT (F) AT (F+EMA) AT (F+CFA)
gaussian_noise 79.0 80.0 79.8 77.6 78.2 76.3 80.2 80.8 78.9
shot_noise 80.0 80.9 80.6 78.5 79.4 77.5 81.3 81.9 80.1
speckle_noise 79.9 80.7 80.5 78.4 79.2 77.3 81.4 81.9 80.0
impulse_noise 74.5 75.1 75.0 73.7 74.5 72.2 75.0 75.2 74.0
defocus_blur 78.6 79.0 78.9 77.9 78.2 76.6 79.9 80.1 77.6
gaussian_blur 76.6 76.8 76.6 76.1 76.3 74.5 77.7 77.6 75.0
motion_blur 74.7 74.9 75.0 74.7 75.0 73.1 75.9 76.0 72.8
zoom_blur 77.6 77.9 77.8 77.2 77.5 75.7 79.4 79.3 76.5
snow 75.6 77.3 77.8 76.3 76.2 77.5 76.7 78.3 77.5
fog 58.1 56.3 55.8 57.8 58.0 53.0 59.1 57.5 52.8
brightness 78.4 79.7 80.1 78.3 78.1 78.2 79.7 80.9 79.8
contrast 42.7 40.5 40.2 41.4 42.0 37.3 43.8 41.7 37.3
elastic_transform 77.4 78.1 78.0 76.9 77.2 75.6 79.1 79.4 77.0
pixelate 81.2 82.0 82.0 80.3 80.6 79.6 82.7 83.1 81.3
jpeg_compression 81.1 81.8 81.8 80.1 80.3 79.6 82.6 83.0 81.4
spatter 77.9 79.1 79.1 78.9 78.3 76.8 79.2 80.0 79.2
saturate 78.7 79.9 79.9 78.1 78.4 77.6 80.3 81.0 80.4
frost 69.3 71.6 72.5 70.7 70.2 70.8 70.9 73.3 70.7
B.2 Overall accuracy results on common corruptions.
We evaluate the robustness and fairness of the adversarial model trained using our approach on common corruptions, and compare it to baseline methods. For this evaluation, we utilize the CIFAR-10C dataset, which contains 18 types of corruptions applied to CIFAR-10 images, using the PreActResNet-18 (PRN-18) architecture.
Table 9 and Table 10 present the overall accuracies for our method and the baselines, respectively. Table 9 shows the overall accuracies of the model using FAIR-TAT, adapted to various approaches on CIFAR-10C, while Table 10 lists the overall accuracies of the baseline methods. The best results on common corruptions for both our method and the baselines are underlined in the tables.
Our approach, FAIR-TAT, particularly when adapted to the baselines, notably FAT, consistently outperforms the baselines in terms of overall model performance on common corruptions. In summary, our findings indicate that FAIR-TAT enhances both robustness and fairness when compared to existing methods on common corruptions. We conclude that FAIR-TAT achieves a better-balanced trade-off between robustness and fairness, as demonstrated by our evaluations on both adversaries and common corruptions.
B.3 Experimental Results on XCiT-S12 on CIFAR10
Most of the works related to fairness consider CNN based models, as they are easy to train adversarially compared to transformer based models. Therefore the fairness is often validated on CNN based models. To further validate the efficacy of FAIR-TAT, we use XCiT-S12, a transformer based model sourced from robustness benchmark. In Table 11, we compare our results to the model trained using vanilla adversarial training. We train XCiT following the recipe of adversarial fine-tuning as mentioned in [12]. Table 11 shows the PGD evaluations of FAIR-TAT vs baseline adversarial training on CIFAR-10. Evaluations in the table indicate that FAIR-TAT is fair on transformer based models as well.
Table 11: Comparison of PGD evaluations of FAIR-TAT framework using XCiT on CIFAR10 with $\epsilon=8/255$. Our method is marked with ◆.
Method XCiT-S12 (Clean Accuracy) XCiT-S12 (Robust Accuracy)
Overall Worst Class Overall Worst Class
AT $89.6$ $79.8$ 62.1 $30.2$
FAIR-TAT ◆ 92.7 80.3 $61.7$ 31.7